### v2

@@ Abstract
Electrocardiogram (ECG) and Phonocardiogram (PCG) signals are linked by a latent coupling signal representing the electrical-to-mechanical cardiac transformation. While valuable for cardiovascular disease (CVD) detection, this coupling signal is traditionally estimated using deconvolution methods that amplify noise, limiting clinical utility. In this paper, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates the problem as distribution matching via optimal transport theory. By jointly optimizing amplitude and temporal alignment, NMCSE mitigates noise amplification without additional preprocessing. Integrated with our Temporal-Spatial Feature Extraction network, NMCSE enables robust multi-modal CVD detection. Experiments on the PhysioNet 2016 dataset with realistic hospital noise demonstrate that NMCSE reduces estimation errors by approximately 30% in Mean Squared Error while maintaining higher Pearson Correlation Coefficients across all tested signal-to-noise ratios. Our approach achieves 97.38% accuracy and 0.98 AUC in CVD detection, outperforming state-of-the-art methods and demonstrating robust performance for real-world clinical applications.
Keywords:
Electrocardiogram Phonocardiogram Coupling Signal Cardiovascular Disease DetectionOptimal Transport
1 Introduction
@@
Cardiovascular disease (CVD), the leading global cause of morbidity and mortality [2], can be effectively managed through early detection and precise diagnosis [10]. Electrocardiogram (ECG) and Phonocardiogram (PCG) are widely utilized non-invasive diagnostic modalities, with ECG providing insights into cardiac electrical activity [5] and PCG capturing heart sounds indicative of mechanical function [20]. However, relying on a single modality often leads to incomplete diagnostic information, thereby increasing the risk of misdiagnosis [15, 9, 1]. As a consequence, multi-modal approaches that integrate ECG and PCG have been developed, effectively combining their complementary features to enhance diagnostic accuracy and reliability [18, 13, 11].
Figure 1: Illustration of the relationship between ECG, PCG, and the coupling signal. ECG, PCG, and the coupling signal serve as three distinct yet complementary modalities for CVD detection. ECG captures the heart’s electrical activity, while PCG reflects its mechanical response. The coupling signal characterizes the transformation from electrical excitation to mechanical contraction, and PCG can be interpreted as the convolution of ECG and the coupling signal. By leveraging these interrelated modalities, the multi-modal approach provides a more comprehensive representation of cardiac function.
The relationship between ECG and PCG is governed by the electromechanical coupling of the heart [23]. The electrical signals (ECG) initiate mechanical contractions that produce the heart sounds (PCG). This relationship can be modeled as PCG being the convolution of ECG with a latent coupling signal that characterizes the electrical-to-mechanical transformation [7]. Incorporating this coupling signal as a third modality can enhance diagnostic performance by explicitly modeling cross-modal dynamics [22].
@@
However, conventional coupling signal estimation methods rely on deconvolution [22, 21], an ill-posed process that amplifies noise in clinical environments. Environmental noise during PCG acquisition renders these approaches unreliable in real-world settings, limiting their practical utility.
To address this challenge, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates the problem as distribution matching via optimal transport theory. This approach: (1) operates at the distribution level to incorporate global signal characteristics; (2) circumvents ill-posed inverse operations; and (3) jointly optimizes amplitude and temporal alignment. We integrate NMCSE with our Temporal-Spatial Feature Extraction network for robust multi-modal CVD detection.
Experiments on the PhysioNet 2016 dataset with simulated hospital noise demonstrate that NMCSE reduces estimation errors by approximately 30% compared to conventional methods and maintains higher correlation across all tested signal-to-noise ratios. Our approach achieves 97.38% accuracy and 0.98 AUC in CVD detection, outperforming state-of-the-art methods in noisy conditions.
2 Related Work
CVD detection has progressed from single-modality to multi-modal approaches that integrate ECG and PCG signals, evolving through three main stages:
@@ 2 Related Work
Feature-level fusion methods extract features independently before combination. Li et al. [18] used specialized networks with genetic algorithm optimization, while Hettiarachchi et al. [12] employed transfer learning on scalograms.
Decision-level fusion methods, as proposed by Li et al. [17], use evidence theory to weight contributions from each modality based on confidence levels.
End-to-end deep learning approaches perform joint representation learning, with Li et al. [16] utilizing a multi-input CNN framework and Zhang et al. [24] developing CPDNet with modality-specific encoders.
These approaches largely treat ECG and PCG as independent information sources. Sun et al. [22] made a significant advance by incorporating coupling signals that explicitly model the electromechanical relationship between ECG and PCG, demonstrating improved diagnostic accuracy. However, their deconvolution-based estimation method is limited by noise amplification, compromising reliability in realistic clinical settings.
Our NMCSE method addresses this critical limitation by providing noise-robust coupling signal estimation, enabling reliable multi-modal CVD detection even in noisy clinical environments.
@@
3 Noise-Robust Multi-Modal Coupling Signal Estimation: Theory and Algorithm
@@
This section establishes the mathematical foundation for applying optimal transport theory to ECG-PCG coupling signal estimation. We develop a comprehensive theoretical framework that addresses the limitations of conventional approaches and provides rigorous error analysis with convergence guarantees.
3.1 Mathematical Model of Cardiac Electromechanical Coupling
We begin by formalizing the relationship between cardiac electrical and mechanical activities within a mathematically rigorous framework.
Theorem 3.1 (ECG-PCG Transformation Model)
@@ 3.1 Mathematical Model of Cardiac Electromechanical Coupling
3.2 Limitations of Deconvolution in Noisy Environments
Theorem 3.2 (Error Amplification in Deconvolution)
@@ 3.2 Limitations of Deconvolution in Noisy Environments
[…]
@@
[…]
Proof (Proof Sketch)
From the frequency domain representation of the noisy PCG signal $\tilde{X}_{\mathrm{PCG}}(\omega)=X_{\mathrm{ECG}}(\omega)H_{t}(\omega)+E(\omega)$, the deconvolution estimate becomes $H_{a}(\omega)=H_{t}(\omega)+\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$. The error term $\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$ is amplified when $X_{\mathrm{ECG}}(\omega)$ approaches zero, resulting in the error bound above.
@@
3.3 Distribution-Based Formulation via Optimal Transport
@@
For this application, the Wasserstein distance from optimal transport theory is particularly well-suited due to three key advantages:
[…]
First, it naturally accounts for the underlying metric structure of the signal space, incorporating both amplitude and temporal aspects essential for preserving physiologically meaningful relationships between cardiac electrical and mechanical events. Second, unlike divergence measures such as KL-divergence, Wasserstein distance exhibits bounded sensitivity to outliers ($W_{p}(P_{1},P_{2})\leq\text{diam}(\text{supp}(P_{1})\cup\text{supp}(P_{2}))$), making it robust to sporadic noise commonly present in clinical recordings.
Most importantly, this approach addresses deconvolution’s inherent instability. The resulting error bounds are independent of the problematic $\frac{1}{X_{\mathrm{ECG}}(\omega)}$ term that causes noise amplification:
@@
This formulation transforms an ill-posed inverse problem into a well-posed optimization problem that remains stable even at frequencies where ECG power approaches zero, with error bounds dependent only on noise properties rather than ECG spectral characteristics. By operating at the distribution level rather than sample-by-sample, our approach inherently incorporates global signal characteristics and statistical properties, making it substantially less susceptible to localized noise artifacts.
3.4 Regularized Optimal Transport and Optimal Cost Function
While the Wasserstein distance provides an elegant theoretical framework, its computational complexity of $O(N^{3})$ can be prohibitive for real-time clinical applications. To address this challenge, we implement a regularized version of optimal transport using the Sinkhorn algorithm, combined with a carefully designed cost function that balances amplitude and temporal fidelity.
Theorem 3.3 (Sinkhorn Distance Properties)
@@ 3.4 Regularized Optimal Transport and Optimal Cost Function
where $H(\Pi)=-\sum_{i,j}\Pi_{i,j}\log\Pi_{i,j}$ is the entropy of the transport plan, achieves computational efficiency, differentiability, and convergence guarantees necessary for coupling signal estimation.
Proof (Proof Sketch)
@@
For cardiac signals with normalized amplitude, we derive an optimal regularization parameter $\varepsilon_{\text{opt}}\approx\frac{\mathbb{E}[C_{i,j}]}{10}\approx 0.1$, balancing approximation accuracy with numerical stability. This specific value ensures the kernel matrix $K_{i,j}=\exp(-C_{i,j}/\varepsilon)$ remains well-conditioned for efficient computation while preserving the method’s noise robustness.
Theorem 3.4 (Optimal Cost Function Design)
@@
Proof (Proof Sketch)
In cardiac electromechanical coupling, both amplitude (contraction strength) and temporal (event timing) characteristics are physiologically significant. For optimal balance, the contribution of each term should be inversely proportional to its expected error variance. For squared error measures ($p=q=2$), this yields the optimal parameter ratio shown above, providing a theoretical foundation for balancing amplitude fidelity and temporal alignment in our cost function.
3.5 NMCSE Algorithm Implementation
@@
3.6 Convergence Properties
Theorem 3.5 (NMCSE Convergence)
@@ 3.6 Convergence Properties
Proof (Proof Sketch)
The Sinkhorn distance is differentiable with respect to input distributions when $\varepsilon>0$, with the gradient with respect to the cost matrix given by $\nabla_{C_{i,j}}\mathcal{D}_{\varepsilon}=\Pi_{i,j}$. By the chain rule and standard SGD convergence assumptions, the algorithm converges to a local minimum at the stated rate. The optimization landscape for cardiac signals exhibits favorable properties with shallow local minima, conferring robustness to initialization.
3.7 Theoretical Comparison with Alternative Methods
Theorem 3.6 (Error Advantage)
@@ 3.7 Theoretical Comparison with Alternative Methods
[…]
@@
4 Temporal-Spatial Feature Extraction Network for ECG, PCG, and Coupling Signals
@@
4.1 Temporal-Spatial Feature Extraction Block
Figure 2: Temporal-Spatial Feature Extraction Block…
@@ 4.1 Temporal-Spatial Feature Extraction Block
4.2 Multi-Modal Processing Pipeline
Figure 3: Multi-Modal Processing Pipeline. Left: Signal spectrogram and its time slices serving as inputs to the spatial and temporal feature extraction branches respectively. Right: System architecture showing ECG, PCG, and coupling signal processing through TSFE modules, multi-modal fusion, and classification.
Table 1 presents the detailed configuration of our proposed Temporal-Spatial Feature Extraction Network.
Table 1: Layer Configuration of the Temporal-Spatial Feature Extraction Network
@@ 4.2 Multi-Modal Processing Pipeline
  • The table outlines the architecture of the proposed TSFE Network.
    • K = Kernel size, $K3\equiv 3\times 3$ for 2D convolutions.
    • S = Stride, controls the step size during convolution.
    • C = Number of output channels at each layer.
    • U = LSTM hidden units, progressively increased across blocks.
    • Each convolutional layer is followed by Batch Normalization (BN) and ReLU.
5 Experiments and Results
This section presents our experimental evaluation of the proposed NMCSE method and TSFE Network. We first assess the robustness of NMCSE for coupling signal estimation under varying noise levels, comparing it with conventional deconvolution methods. We then evaluate the effectiveness of integrating the NMCSE-estimated coupling signal with ECG and PCG as multi-modal inputs to the TSFE network for CVD detection.
5.1 Datasets and Data Preprocessing
5.1.1 PhysioNet/CinC Challenge 2016 Dataset
We utilized the PhysioNet/CinC Challenge 2016 dataset [6, 19], specifically subset A containing 405 synchronized ECG-PCG recordings from 121 subjects. Signals were resampled to 2000 Hz with durations ranging from 9 to 37 seconds. To our knowledge, this represents the only publicly available dataset with synchronized ECG-PCG signals suitable for CVD detection. All experiments employed five-fold cross-validation.
5.1.2 Hospital Ambient Noise Dataset
To simulate clinical environments, we augmented the PCG signals with noise from the Hospital Ambient Noise Dataset [3], which includes common hospital sounds such as equipment noise, staff conversations, patient movements, and ambient environmental sounds. Each PCG recording was mixed with randomly selected noise at various signal-to-noise ratios (SNRs) through linear additive noise augmentation to evaluate coupling signal estimation robustness under conditions typically encountered in clinical settings.
5.1.3 Data Preprocessing

### v3

@@ Abstract
The coupling signal refers to a latent physiological signal that characterizes the transformation from cardiac electrical excitation, captured by the electrocardiogram (ECG), to mechanical contraction, recorded by the phonocardiogram (PCG). By encoding the temporal and functional interplay between electrophysiological and hemodynamic events, it serves as an intrinsic link between modalities and offers a unified representation of cardiac function, with strong potential to enhance multi-modal cardiovascular disease (CVD) detection. However, existing coupling signal estimation methods remain highly vulnerable to noise, particularly in real-world clinical and physiological settings, which undermines their robustness and limits practical value. In this study, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates coupling signal estimation as a distribution matching problem solved via optimal transport. By jointly aligning amplitude and timing, NMCSE avoids noise amplification and enables stable signal estimation. When integrated into a Temporal-Spatial Feature Extraction (TSFE) network, the estimated coupling signal effectively enhances multi-modal fusion for more accurate CVD detection. To evaluate robustness under real-world conditions, we design two complementary experiments targeting distinct sources of noise. The first uses the PhysioNet 2016 dataset with simulated hospital noise to assess the resilience of NMCSE to clinical interference. The second leverages the EPHNOGRAM dataset with motion-induced physiological noise to evaluate intra-state estimation stability across activity levels. Experimental results show that NMCSE consistently outperforms existing methods under both clinical and physiological noise, highlighting it as a noise-robust estimation approach that enables reliable multi-modal cardiac detection in real-world conditions.
I Introduction
@@
Cardiovascular disease (CVD), the leading global cause of morbidity and mortality [1], can be effectively managed through early detection and precise diagnosis [2]. Electrocardiogram (ECG) and Phonocardiogram (PCG) are widely utilized non-invasive diagnostic modalities, with ECG providing insights into cardiac electrical activity [3] and PCG capturing heart sounds indicative of mechanical function [4]. However, relying on a single modality often leads to incomplete diagnostic information, thereby increasing the risk of misdiagnosis [5]. As a consequence, multi-modal approaches that integrate ECG and PCG have been developed, effectively combining their complementary features to enhance diagnostic accuracy and reliability [6, 7, 8].
Figure 1: Illustration of the relationship between ECG, PCG, and the coupling signal. ECG reflects the heart’s electrical activity, while PCG conveys its mechanical response. The coupling signal captures the dynamic transformation between them.
The relationship between ECG and PCG originates from the electromechanical coupling of the heart, where electrical excitation governs mechanical contraction [9]. Reflecting this physiological process, prior studies have modeled the PCG as the output of a system driven by the ECG, often formulated as a convolution with a latent impulse response, known as the coupling signal [10]. This signal captures the temporal mapping from electrical to mechanical domains, reflecting how electrical activity propagates through the myocardium to produce acoustic responses. Estimating the coupling signal enables explicit modeling of cross-modal dynamics and has been shown to improve diagnostic performance when incorporated as a third modality alongside ECG and PCG [11].
@@
However, existing methods for coupling signal estimation are typically based on deconvolution [11, 12], an ill-posed inverse operation that is highly sensitive to noise [13]. In practice, environmental interference during PCG acquisition—especially in clinical settings—renders such approaches unstable and limits the utility in real-world applications.
To address these limitations, we propose Noise-Robust Multi-Modal Coupling Signal Estimation (NMCSE), which reformulates the estimation task as a distribution matching problem solved via optimal transport. Rather than recovering the coupling signal through point-wise inversion, NMCSE aligns the global distributions of ECG-derived and PCG signals, allowing for more stable estimation under noise. We further introduce a Temporal-Spatial Feature Extraction (TSFE) network to integrate the estimated coupling signal with ECG and PCG for multi-modal classification, where convolutional and recurrent modules are jointly used to capture both spectral patterns and temporal dependencies.
To evaluate the robustness of the proposed method, we conduct two complementary experiments: one on the PhysioNet 2016 dataset [14, 15] with hospital noise [16], assessing resilience to clinical interference; and another on the EPHNOGRAM dataset [17], evaluating stability under real motion-induced physiological noise. In both settings, NMCSE consistently outperforms baseline methods in terms of accuracy, correlation, and estimation consistency—achieving 97.38% classification accuracy and 0.98 AUC—demonstrating its practical utility in multi-modal cardiac analysis.
II Related Work
II-A Multi-Modal Fusion Methods for CVD Detection
Recent advancements in cardiovascular disease (CVD) detection have transitioned from single-modality approaches to multi-modal frameworks that integrate electrocardiogram (ECG) and phonocardiogram (PCG) signals. These efforts have primarily followed three fusion paradigms:
@@ II-A Multi-Modal Fusion Methods for CVD Detection
Feature-level fusion methods extract features from each modality independently and then concatenate or combine them for classification. For instance, Li et al. [6] designed modality-specific neural networks with genetic algorithm-based optimization, while Hettiarachchi et al. [18] applied transfer learning on PCG scalograms combined with ECG features.
Decision-level fusion methods produce individual predictions from separate models and combine them using ensemble strategies. Li et al. [19] applied evidence theory to weight the decisions of each modality according to their estimation.
End-to-end deep learning approaches enable joint representation learning from raw signals. Li et al. [20] proposed a multi-input CNN to learn from both ECG and PCG, while Zhang et al. [21] developed CPDNet, which utilizes modality-specific encoders and cross-modal attention mechanisms.
Despite their differences, these approaches treat ECG and PCG as independent or weakly correlated inputs, without explicitly modeling the physiological relationship between electrical and mechanical cardiac activity. However, recent studies suggest that explicitly modeling this relationship can enhance multi-modal performance, indicating the importance of incorporating physiological coupling into fusion frameworks—a direction we further elaborate in the next subsection.
II-B Transformer-Based Multi-Modal Learning Methods
Recent studies have introduced Transformer architectures into multi-modal cardiovascular signal analysis to enhance feature fusion between ECG and PCG. For example, CAD‑ViT [22] employed a dual-modal Vision Transformer with co-attention mechanisms and dynamic weighted fusion to model cross-modal dependencies between electrical and mechanical signals. PACFNet [23] proposed a progressive attention-based framework, stacking multiple self-attention layers to integrate modality-specific features. However, these approaches often depend on extensive labeled data and exhibit vulnerability to random noise, limiting their applicability in real-world clinical scenarios.
II-C Coupling Signal Estimation Methods
To incorporate the physiological relationship between ECG and PCG into multi-modal frameworks, a common approach is to estimate the latent coupling signal that characterizes the electromechanical transformation of the heart. This signal is typically modeled as the impulse response of a linear time-invariant system, with the PCG viewed as the convolution of the ECG and the coupling filter.
Several studies have adopted deconvolution-based techniques to estimate this signal. Representative methods include classical inverse filtering, Tikhonov-regularized deconvolution, Wiener filtering with estimated noise spectra, and sparsity-constrained recovery. Sun et al. [11] demonstrated that such estimated coupling signals, when used as an auxiliary modality, can improve CVD detection performance.
Despite their utility, these methods are fundamentally ill-posed and prone to instability in noisy conditions. When the ECG exhibits low spectral energy at certain frequencies, spectral division during deconvolution can amplify noise and lead to unreliable estimates—especially in real-world clinical or ambulatory environments [24].
This limitation has prompted exploration of alternative formulations that focus on distribution-level alignment rather than point-wise inversion. Building on this idea, we propose a noise-robust multi-modal coupling signal estimation method (NMCSE) based on optimal transport, which aligns the global distributions of ECG-transformed and PCG signals. This approach avoids noise amplification and enables stable coupling signal estimation even under real-world clinical noise.
@@ II-C Coupling Signal Estimation Methods
III Noise-Robust Multi-Modal Coupling Signal Estimation: Theory and Algorithm
@@
This section establishes the mathematical foundation for applying optimal transport to ECG-PCG coupling signal estimation. We develop a comprehensive theoretical framework that addresses the limitations of deconventional approaches and provides rigorous error analysis with convergence guarantees.
III-A Model of Cardiac Electromechanical Coupling
Theorem 1 (ECG-PCG Transformation Model).
@@ III-A Model of Cardiac Electromechanical Coupling
III-B Limitations of Deconvolution in Noisy Environments
Theorem 2 (Error Amplification in Deconvolution).
@@ III-B Limitations of Deconvolution in Noisy Environments
$\|H_{a}-H_{t}\|_{2}^{2}=\int_{-\infty}^{+\infty}\left|\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}\right|^{2}d\omega$ (3)
@@
$\|H_{a}-H_{t}\|_{2}^{2}>\frac{1}{\delta^{2}}\int_{\Omega_{\delta}}|E(\omega)|^{2}d\omega$ (4)
Proof Sketch.
From the frequency domain representation of the noisy PCG signal $\tilde{X}_{\mathrm{PCG}}(\omega)=X_{\mathrm{ECG}}(\omega)H_{t}(\omega)+E(\omega)$, the deconvolution estimate becomes $H_{a}(\omega)=H_{t}(\omega)+\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$. The error term $\frac{E(\omega)}{X_{\mathrm{ECG}}(\omega)}$ is amplified when $X_{\mathrm{ECG}}(\omega)$ approaches zero, resulting in the error bound above. ∎
@@
III-C Distribution-Based Formulation via Optimal Transport
@@
The Wasserstein distance from optimal transport theory is particularly well-suited due to three key advantages:
$W_{p}(P_{1},P_{2})=\left(\inf_{\gamma\in\Gamma(P_{1},P_{2})}\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}d\gamma(x,y)\right)^{1/p}$ (6)
First, it naturally accounts for the underlying metric structure of the signal space, incorporating both amplitude and temporal aspects essential for preserving physiologically meaningful relationships between cardiac electrical and mechanical events. Second, unlike divergence measures such as KL-divergence, Wasserstein distance exhibits bounded sensitivity to outliers ($W_{p}(P_{1},P_{2})\leq\text{diam}(\text{supp}(P_{1})\cup\text{supp}(P_{2}))$), making it robust to sporadic noise present in clinical recordings.
Most importantly, our approach overcomes the instability of deconvolution. Unlike existing methods whose error grows with $\frac{1}{X_{\mathrm{ECG}}(\omega)}$, our error bound depends only on noise:
@@
This reformulation turns an ill-posed inversion into a well-posed optimization, ensuring stability even when ECG power is low. By aligning distributions rather than individual samples, it captures global signal structure and resists noise artifacts.
III-D Regularized Optimal Transport and Optimal Cost Function
While the Wasserstein distance provides a theoretical framework, its computational complexity of $O(N^{3})$ can be prohibitive for clinical applications. To address this challenge, we implement a regularized version of optimal transport using the Sinkhorn algorithm, combined with a carefully designed cost function that balances amplitude and temporal fidelity.
Theorem 3 (Sinkhorn Distance Properties).
@@ III-D Regularized Optimal Transport and Optimal Cost Function
where $H(\Pi)=-\sum_{i,j}\Pi_{i,j}\log\Pi_{i,j}$ is the entropy of the transport plan, achieves computational efficiency and convergence guarantees necessary for coupling signal estimation.
Proof Sketch.
@@
For cardiac signals with normalized amplitude, we derive an optimal regularization parameter $\varepsilon_{\text{opt}}\approx\frac{\mathbb{E}[C_{i,j}]}{10}\approx 0.1$, balancing approximation accuracy with numerical stability. This specific value ensures the kernel matrix $K_{i,j}=\exp(-C_{i,j}/\varepsilon)$ remains well-conditioned for efficient computation while preserving the method’s noise robustness. ∎
Theorem 4 (Optimal Cost Function).
@@
Proof Sketch.
In cardiac electromechanical coupling, both amplitude and timing are physiologically important. To balance them optimally, each term’s contribution should be inversely proportional to its error variance. For squared error measures ($p=q=2$), this yields the optimal parameter ratio, providing a theoretical basis for our cost function.
III-E NMCSE Algorithm Implementation
@@
III-F Convergence Properties
Theorem 5 (NMCSE Convergence).
@@ III-F Convergence Properties
Proof Sketch.
The Sinkhorn distance is differentiable with respect to input distributions when $\varepsilon>0$, with the gradient with respect to the cost matrix given by $\nabla_{C_{i,j}}\mathcal{D}_{\varepsilon}=\Pi_{i,j}$. By the chain rule and standard SGD convergence assumptions, the algorithm converges to a local minimum at the stated rate. ∎
III-G Theoretical Comparison with Alternative Methods
Theorem 6 (Error Advantage).
@@ III-G Theoretical Comparison with Alternative Methods
$\int_{\Omega_{\delta}}\frac{|E(\omega)|^{2}}{f(|X_{\mathrm{ECG}}(\omega)|)}d\omega$ (13)
@@
IV Temporal-Spatial Feature Extraction Network for ECG, PCG, and Coupling Signals
@@
IV-A Temporal-Spatial Feature Extraction Block
Figure 2: Temporal-Spatial Feature Extraction Block.
@@ IV-A Temporal-Spatial Feature Extraction Block
TABLE I: Layer Configuration of the Temporal-Spatial Feature Extraction Network
@@
Notation: K = kernel size ($K3\equiv 3\times 3$ for 2D convolutions); S = stride; C = number of output channels; U = number of LSTM hidden units (increasing by block); All convolutional layers are followed by Batch Normalization and ReLU.
IV-B Multi-Modal Processing model
Figure 3: Multi-Modal Processing Pipeline. Left: Signal spectrogram and time slices as inputs to spatial and temporal branches. Right: ECG, PCG, and coupling signals processed via TSFE, fused, and classified.
Table I presents the detailed configuration of our proposed Temporal-Spatial Feature Extraction Network.
V Experiments and Results
V-A Datasets and Data Preprocessing
V-A1 PhysioNet/CinC 2016
We used the PhysioNet/CinC Challenge 2016 dataset [14, 15], specifically subset A containing 405 synchronized ECG-PCG recordings from 121 subjects. All signals were resampled to 2000 Hz, with durations ranging from 9 to 37 seconds. To the best of our knowledge, this is the only publicly available dataset that provides synchronized ECG and PCG recordings suitable for CVD detection. Five-fold cross-validation was applied throughout.
V-A2 Hospital Ambient Noise
To simulate clinical noise conditions, we augmented the PCG recordings with sounds from the Hospital Ambient Noise Dataset [16], which includes equipment beeps and ambient hospital sounds. Noise was mixed at various signal-to-noise ratios (SNRs) using linear additive augmentation to evaluate the robustness of coupling signal estimation under external interference.
V-A3 EPHNOGRAM
The EPHNOGRAM dataset [17] contains synchronized ECG and PCG recordings collected across multiple physical activities, including rest, walking, and stair climbing. It reflects realistic physiological variability and motion-induced noise. We used this dataset to assess the consistency of coupling signal estimation within the same physiological state, providing a complementary view of robustness beyond artificial noise injection.
V-A4 Data Preprocessing